Nightly Observing Plan for MJD 56631 (starting evening of Dec 4, 2013)


Observer 1 Observer 2 Run Manager Assistant Run Manager SISPI Expert on Call
Rutu Das Elisabeth Krause Lyndsay Old Marcelle Soares-Santos
ext. 521 ext. 515 ext. 516 ext. 518

CTIO staff (Telescope Operators)

Day Support Night Assistant
Manuel Hernandez Claudio Aguilera

DECam Configuration and Setup

SISPI Configuration File DES.ini
Proposal ID 2012B-0001
Image Processing (Calibration) NCSA
Image Processing (First Cut) NCSA

mjd 56631, date 20131204, Site CTIO, UT-offset 3 hr
Current local time 10:31 CLST     UT 13:31     SidTim 13:43
Sunset     20:41 CLST  23:41 UT   SidTim  358 deg  
Eve. -10   21:21 CLST  00:21 UT   SidTim    8 deg  Zeropoint; standards
Eve. -14   21:44 CLST  00:44 UT   SidTim   14 deg  Begin obstac
Eve. -18   22:08 CLST  01:08 UT   SidTim   20 deg  Last chance standards
Moonset    22:18 CLST  01:18 UT   SidTim   23 deg  
Midpoint   01:33 CLST  04:33 UT   SidTim   72 deg  
Morn -18   04:59 CLST  07:59 UT   SidTim  123 deg  Begin astronomical twilight
Morn Std   05:23 CLST  08:23 UT   SidTim  129 deg  End obstac; begin standards
Morn -10   05:45 CLST  08:45 UT   SidTim  134 deg  
Sunrise    06:26 CLST  09:26 UT   SidTim  144 deg  
Moon position (deg): RA 283.8, Dec -17.3 Phase: 0.160

Weather Conditions

*Tololo Weather Links .

Observing Plan and Goals for the Night

  1. Afternoon calibrations script to use: prenight_calibration.json
  2. Start of night standards (preceed with zeros script: script to use: flush_Vsub.json).
  3. Nightly out of focus image. script to use: defocus_nightly.json (right after last standard, no break)
  4. ObsTac.
  5. End of night standards.
  6. End of night calibrations. if start of night calibrations good use: postnight_calibration.json, otherwise use postnight_zeros_flats.json

Start of Night Standards MANUAL SLEW TO THESE FIELDS Wind screen down.
UT Date = '2013/12/05 00:21' Local time = 21:21
E8-A ra 20:07:22 dec -44:37:01 secz 1.7 script to use: DES_std_e8a_Yzirg.json
SDSSJ2140-0000 ra 21:40:00 dec +00:00:00 secz 1.6 script to use: DES_std_sdssj2140_Yzirg.json
SDSSJ0100-0000 ra 01:00:00 dec +00:00:00 secz 1.2 script to use: DES_std_sdssj0100_Yzirg.json

End of Night Standards MANUAL SLEW TO THESE FIELDS Check Wind screen down.
UT Date = '2013/12/05 08:23' Local time = 05:23
E2-A ra 04:03:00 dec -44:41:45 secz 1.7 script to use: DES_std_e2a_grizY.json
SDSSJ1048-0000 ra 10:48:00 dec +00:00:00 secz 1.4 script to use: DES_std_sdssj1048_grizY.json
SDSSJ0843-0000 ra 08:43:00 dec +00:00:00 secz 1.2 script to use: DES_std_sdssj0843_grizY.json

Notes from 4pm Meeting

Run Manager's Meeting Dec. 04, 2013
Brian N., HTD, Eric, Lyndsay, Marcelle, Manuel
Josh, Margaret, Robert

Telescope and Camera (Manuel)
Last night there was a problem with the Dome Encoder. It was fixed today.
People working on the new Dome cooling system.
Also, pumping on on the LN2 lines.

Telescope and Camera will be ready to go about 4:30. SISPI
needs to be restarted.

Last Night (Lyndsay)
Obstac all night, modest seeing. WF only. Lots of occluded images.
SK detected a lot of clouds during the 2nd 2/3 of the night. We
wonder if there is a correlation between this and the occluded images.
The clouds last night came in during the 2nd half.

A list of definitely occluded images is in the night summary.
Every time there was a SISPI alarm for delta (dome - tel) they
looked at donuts and could see by eye that there was a problem.

12/1 DQ is done.

In Re last night's occluded images: these will be declared bad by DESDM (unprocessable).
DESDM will send a CSV to the shift.

weather is the same as last night.
Obstac: tiling 3, maybe 4, and some possibility of SN because some
fields will be a latency of 4.

Everybody should have bright flashlights. If NG, see telops.